1. Find the specific function values f(x, y) = x^2 + xy^3


(a) f(2, 3) (b) f(−3, −2)


1
Expert's answer
2022-03-19T04:13:40-0400

"f\left( 2,3 \right) =2^2+2\cdot 3^3=58\\f\left( -3,-2 \right) =\left( -3 \right) ^2+\left( -3 \right) \cdot \left( -2 \right) ^3=33"